can you please help solve a polynomial? x^2-34x+325
Answer by stanbon(75887)

solve a polynomial? x^2-34x+325
----------------------------------
a = 1 ; b = -34 ; c = 325
------------------------------
Quadratic Formula:
x = [34 +- (34^2 - 4*1*325)]/2
------
x = [34 +- sqrt(-144)]/2
---
x = [34 +- 12i]/2
---
x = [17 +- 6i]
--------------------
x = 17 + 6i or x = 17-6i
============================
Cheers,
Stan H.